In a music award ceremony,5 female singers are competing for gold,silver,and bronze awards under the best female singers
If kate is one of the winners,how many permutations are there for the winners in the 3 awards?

Answer by ikleyn(52778)

.

If Kate won gold, then for the rest candidates there are 4*3 = 12 possible permutations.


If Kate won silver, then for the rest candidates there are 4*3 = 12 other possible permutations.


If Kate won bronze, then for the rest candidates there are 4*3 = 12 possible other permutations.


These three sets of permutations with 12 elements in each set, are DISGIUNT, 

so the final answer is  12+12+12 = 36 permutations.   

ANSWER. In all, there are 36 possible permutations.


Solved.


.............

Now you see this advantage posting separate problems/questions.